### v2

# Can Large Language Models Capture Video Game Engagement?

Abstract
Can out-of-the-box pretrained Large Language Models (LLMs) detect human affect successfully when observing a video? To address this question, for the first time, we evaluate comprehensively the capacity of popular LLMs for successfully predicting continuous affect annotations of videos when prompted by a sequence of text and video frames in a multimodal fashion. In this paper, we test LLMs’ ability to correctly label changes of in-game engagement in 80 minutes of annotated videogame footage from 20 first-person shooter games of the GameVibe corpus. We run over 4,800 experiments to investigate the impact of LLM architecture, model size, input modality, prompting strategy, and ground truth processing method on engagement prediction. Our findings suggest that while LLMs rightfully claim human-like performance across multiple domains and able to outperform traditional machine learning baselines, they generally fall behind continuous experience annotations provided by humans. We examine some of the underlying causes for a fluctuating performance across games, highlight the cases where LLMs exceed expectations, and draw a roadmap for the further exploration of automated emotion labelling via LLMs.
Index Terms:
Large language models, affective computing, player modelling, engagement
I Introduction
The use of autoregressive modelling and large pretrained models such as Large Language Models (LLMs) is currently dominating AI research. LLMs have demonstrated unprecedented advances in language translation, code generation, problem solving, and AI-based assistance among many other downstream tasks [59]. Given their versatility and efficiency compared to earlier autoregressive models, one might even argue that the current capabilities of LLMs are endless as long as a problem and its corresponding solution(s) are represented as text. Meanwhile, the recent applications of LLMs within affective computing largely consider text-based affect modelling tasks such as LLM-based sentiment analysis [87, 13, 88]. The automatic labelling of affect based on time-continuous visual input remains largely unexplored [88]. The handful of studies focusing on multimodal input, still largely rely on images [35, 80] or—in rare cases where we see video and/or audio input [92]—they focus on larger units, evaluating whole scenes instead of time-continuous labelling.
Figure 1: Clips in the GameVibe Dataset. List of game titles: (1) Apex Legends; (2) Blitz Brigade; (3) Borderlands 3; (4) Corridor 7; (5) Counter Strike 1.6; (6) CS:GO - Dust2; (7) CS:GO - Office; (8) Doom; (9) Insurgency; (10) Far Cry; (11) Fortnite; (12) Heretic; (13) Medal of Honor 2010; (14) Overwatch 2; (15) PUBG; (16) Medal of Honor 1999; (17) Team Fortress 2; (18) Void Bastards; (19) HROT; (20) Wolfram.
Motivated by the aforementioned lack of studies, this paper introduces the first comprehensive evaluation of LLMs tasked to predict time-continuous affect labels from videos. We observe the ability of various LLMs to rank consecutive video frames in terms of viewer engagement. While engagement is generally not considered a well-defined emotional reaction, emotional engagement encompasses a broad range of measurable affective reactions—including evaluative judgments, facial changes, autonomic responses (such as skin conductance and heart rate) [18]—that strongly correlate with user preferences [7], and can be conceptualized as a high arousal, positive valence affective response under the Circumplex Model of affect [63]. Throughout this paper we use engagement as an affective measure that captures the complexity of a positive emotional experience [7] of multimedia viewing. We chose games as the domain of our study since they can act as rich elicitors of emotions and can offer a wide range of dynamic scenes and stimuli, varying from intense player actions to less intense game-world exploration. Even though LLMs have been used in a series of diverse tasks within the domain of videogames—both in academic studies [26, 84] and industrial applications such as AI Dungeon (Latitude, 2019), AI People (GoodAI, 2025) and Infinite Craft [^1] https://neal.fun/infinite-craft/—the capacity of these foundation models as predictors of player experience has not been investigated yet.
Figure 2: Overview of the evaluation experiments presented in this study. Independently of experimental setting, the downstream task is engagement prediction formulated as a binary preference. We use a combination of text prompts and/or video frames as input and task the LLMs to label engagement. To evaluate the models, we compare the generated labels to the ground truth labels from the annotated GameVibe corpus (see Section III). All LLMs are prompted with a Chain-of-Thought (CoT) strategy. We experiment with three different setups for possible inputs: Multimodal One-Shot—based on image input; Text Descriptions—based on LLM-generated descriptions of single frames; and Multimodal Few-Shot—based on LLM-generated positive and negative examples of engagement evaluation based on the ground truth. In all experimental settings we generate a description, comparison, reasoning, and a decision relating to an increase or decrease in engagement. We parse these outputs to derive the final binary engagement evaluation. Descriptions are only generated in the Multimodal settings.
We employ LLMs as autonomous player experience annotators and present a thorough evaluation of their capacity to predict player experience in one-shot and few-shot fashions. Specifically, we compare state of the art foundation models from the LLaVA and GPT families against human annotated data of viewer engagement of the GameVibe dataset [8] (see Fig. 1). The dataset contains continuous engagement labels of gameplay videos across a variety of first-person shooter (FPS) games. We present selected results out of $4,840$ experimental settings in which we vary and test LLM model types, model sizes, prompting strategies, input types, and ground truth processing methods across $20$ games. Figure 2 shows a high level overview of our experimental setup followed in this study. We focus on videogame footage—one might come across on game streaming services such as Twitch [^2] https://www.twitch.tv—as input and viewer engagement as output.
The novelty of this paper is two-fold. First, we investigate the capacity of LLMs to accurately label affect in a time-continuous manner using videos as affect elicitors. Second, we present the first large set of evaluation experiments that lays the groundwork for LLM-based player experience prediction. Our experiments show the feasibility of leveraging LLMs for engagement prediction, particularly on popular games with a rich online presence (such as ApexLegends, 2019). Our key findings suggest that a) text-based summarisation of frames and direct multimodal prompting do not impact LLM performance; b) LLM performance is largely dependent on the elicitor (i.e. different games in this study); c) the multimodal few-shot prompting strategy is the one that improves LLM performance the most; and d) scale matters. Specifically, the best results obtained are when we employ the GPT-4o model and we feed it with a few positive and negative multimodal examples of increasing or decreasing engagement (few-shot prompting). While this approach yields an average accuracy of $6\%$ over a Zero Rule Classifier across games, the GPT-4o model is able to improve the baseline performance by up to $47\%$ in certain games. Even though this performance is far from consistent, all tested LLMs exceeded traditional machine learning baselines in every experiment, underscoring their promise for complex affect modelling tasks.
The paper is structured as follows. Section II presents related work on LLMs for affect modelling, uses of LLMs in games, and player modelling. Section III briefly presents the GameVibe dataset and the data preprocessing. Section IV discusses our approach, presenting the models used and the different prompting strategies we employed. Section V presents the key results obtained, including a sensitivity analysis and hyperparameter tuning, a comparison between different input modalities, results of few-shot experiments, and a qualitative analysis on the most and least successful models. The paper ends with a brief discussion on possible avenues for future research (see Section VI) and our key conclusions (see Section VII).
II Related Work
This study investigates the capacity of LMMs to accurately annotate subjectively-defined aspects of gameplay. We leverage the existing knowledge-priors of these algorithms, without fine-tuning or using complex retrieval augmented strategies. We thus hypothesise that the algorithm’s prior knowledge is sufficient to approximate the ground truth of engagement (as provided via human feedback) in a set of gameplay scenarios. This section covers related work in affect modelling using LLMs, the use of LLMs in games, and it ends with a focus on modelling aspects of players and their games.
II-A LLMs for Affect Modelling
Given the resounding success of LLMs in several domains, several recent research efforts naturally focus on their direct application in affect detection tasks. The vast majority of research on LLMs related to human affect have focused on predicting manifestations of affect from text as this plays to the strengths of their architecture. Unsurprisingly, sentiment analysis has been the most common research application of LLMs in affective computing and has given us some impressive results already [41]. Indicatively, Broekens et al. [13] highlighted how GPT-3.5 can accurately perform sentiment analysis on the ANET corpus [12] for valence, arousal and dominance. Similarly, Müller et al. [49] used fine-tuned Llama2-7b [70] and Gemma [69] models to classify shame in the DEEP corpus [64], achieving $84\%$ accuracy. Whilst LLMs have been extensively tested for sentiment analysis on existing text-based corpora, research on using LLMs as predictors of experience by observing multimodal content such as games remains unexplored.
Despite their promise, some critical challenges have emerged when working with pre-trained LLMs for prediction tasks such as affect modelling. A recent study by Chochlakis et al. [17] has found that LLMs struggle to perform meaningful in-context learning from new examples and remain fixed to their knowledge priors, with larger models exaggerating this issue. This problem is even more pressing in closed-source models such as GPT-4o because researchers lack important details which can help them assess the level of data contamination. Balloccu et al. [6] conducted a study across $255$ academic papers and found that LLMs have been exposed to a significant number of samples from existing ML benchmarks, potentially painting a misleading picture about their predictive performance in such tasks. While the dataset we use in this paper covers a novel domain, it is possible that some of the videos in the GameVibe dataset have been exposed to some of the models we use. However, because the dataset was published after the models used here [^3] LLaVA 1.6 was published on 18 July 2023; GPT-4o was originally released on 13 May 2024; the GameVibe dataset was published on 17 June 2024., we are confident that the engagement prediction task specifically does not suffer from any significant data contamination.
Beyond contamination, we also have to face the inherent biases encoded in LLMs. Mao et al. [41] have conducted a study on such biases in BERT-like models [22] on affective computing tasks. In our study we use what Mao et al. call “coarse-grain” tasks—a binary decision with symmetrical labels (here increase and decrease of engagement). When evaluating these types of tasks, LLMs have been shown to exhibit less bias [41] than on “fine-grained” tasks with multiple asymmetrical labels. This gives us confidence on the feasibility of our task—which is formulated as a binary classification problem.
Amin et al. [3] have also conducted a study on the capabilities of GPT [51] on affective computing tasks. They have put forth a comprehensive series of experiments which included a similar pairwise preference classification task for engagement prediction to what we use in this paper. They showed that when it comes to subjective tasks with a high potential for disagreement between annotators, out-of-box LLMs, such as GPT struggle compared to architectures leveraging specialized supervised networks. In those experiments—focusing on a simple one-shot prompting strategy on text input—GPT barely surpassed the baseline. In contrast to [3], we investigate multimodal, chain-of-thought, and few-shot strategies in visual-based engagement prediction tasks across multiple games, analysing where LLMs either struggle or flourish compared to baseline approaches.
While our work examines the capabilities of general-purpose LLMs without any task-specific fine-tuning, recent research has increasingly turned to building specialized multimodal architectures and datasets. In particular, HumanOmni [92] introduces a large-scale corpus of human-centric data and an adaptive multimodal fusion architecture fine-tuned for emotion recognition, facial expression description, and action understanding. Similarly, FaVChat [91] focuses on human-centric content, but narrows its scope to fine-grained facial video analysis. While such domain-specific approaches demonstrate clear performance gains, they rely on extensive pre-training over curated datasets. In contrast, here we evaluate general-purpose, out-of-the-box models without additional fine-tuning.
II-B LLMs in Games
The recent developments in LLM methods and technology brought unprecedented wide adoption of AI across multiple domains including law [32], healthcare [50], and education [48]. Advancements in transformer architectures [73], coupled with a rapid increase in dataset and parameter sizes [31], led to a new wave of algorithms with previously unseen capabilities to generate high-quality text. Starting with Bidirectional Encoder Representations from Transformers (BERT) [22] and eventually popularized with the release of Generative Pre-trained Transformers (GPT) [59, 24, 51], LLMs have largely been characterized as transformer-based models, using large amounts of parameters (in the 100 millions and billions), built on large amounts of data, generating text in an autoregressive manner—that is predicting future tokens based on prior data. More recently, LLMs have been expanded to handle new modalities beyond text, such as audio and images [70], making them a candidate for applications using multimodal content such as gameplay videos.
In the context of games, LLMs have been used to create game-playing agents [72, 29], commentators [61] game analytics [75, 62], AI directors and game masters [85, 93], content generators [67], and design assistants [25]. Beyond the academic setting, we are seeing considerable interest from industrial players as well, such as NVIDIA’s recent ACE small language models  [^4] https://developer.nvidia.com/ace for autonomously generating the behaviour and animation of NPCs. Gallotta et al. [26] offer a recent and thorough overview on how LLMs can be utilised in games. In their roadmap, they identify player modelling as one of the most promising, yet unexplored avenues for future research into LLMs and games. Whilst affect modelling research has demonstrated that LLMs can be effective predictors in tasks such as sentiment analysis [41], they are yet to be widely evaluated for modelling player experience in the context of games.
II-C Player Affect Modelling
Player modelling is an active field within AI and games research [84] with a particular focus on methods that capture emotional and behavioural aspects of gameplay. Studies into player affect modelling often focus on arousal and valence [63]—aiming to deconstruct this emotional experience [45]. However, under the same theoretical paradigm, we can also conceptualize more complex emotional patterns as a function of affective dimensions. Given the complexity of both game playing and game spectating [7, 19], multiple studies have focused on higher level constructs, such as fun [9] and engagement [43, 8, 57]. Engagement in both gameplay and spectatorship can be characterized as a high arousal state [30]—with a positive valence in contrast to anxiety or frustration [19].
Traditionally, the field has focused heavily on data aggregation [23] and pattern discovery [40, 42] of playing behaviours, toxicity [14] and motivation [42]; but there has been a recent shift towards moment-to-moment predictive models of players [39, 43, 45, 11, 8]. The prevalent strategy of such modelling methods relies on the availability of continuous annotation traces, which are generally processed as interval data [82]. This allows for the treatment of the labelled data as absolute ratings or classes such as low and high intensity [11, 39].
In contrast to the traditional way of treating annotations as absolute ratings, here we view player modelling as an ordinal learning paradigm aiming to maximize the reliability and validity of our predictive models [82, 81]. We task LLMs to label increases or decreases of engagement across frames of a game instead of asking them to provide ratings of engagement per frame. The ordinal representation of subjective notions such as engagement is supported both by theories of human psychology and cognition [28, 66] and by a growing body of research in neuroscience [20] and affective computing [83, 38, 82, 47, 45, 8] among other disciplines. Importantly, we employ LLMs and we test their ability to model game engagement as viewed through gameplay videos.
Figure 3: Example clip from GameVibe showcasing the annotation interface using PAGAN and the RankTrace annotation tool for collecting unbounded, time continuous signals in real-time.
III The GameVibe Corpus
This section gives a general overview of the GameVibe corpus used throughout all experiments presented in this paper followed by an outline of the preprocessing approach we adopted for the engagement labels in this study. While the dataset is introduced thoroughly in [8], in this section we highlight the main aspects of the dataset that are relevant to our experiments here.
TABLE I: Core properties of the original Gamevibe corpus and the processed version (GameVibe-LLM) used in this study
Properties GameVibe GameVibe-LLM
Annotators 20 20
Number of videos 120 videos 80 videos
Video database size 120 minutes 80 minutes
Number of games 30 games 20 games
Gameplay video duration 1 minute each 1 minute each
Annotation type Interval signal Discrete ordinal
Modalities Visual, audio Visual
III-A Corpus Overview
The GameVibe corpus [8] consists of a set of $120$ audiovisual clips and human annotations for engagement as viewers of first-person shooter games. This corpus presents a significant challenge for affect modelling research as its stimuli encompass a wide variety of graphical styles (e.g. photorealistic, retro) and game modes (e.g. deathmatch, battle royale). Table I contains a basic summary of the properties of this corpus and processed version we use for this study.
Each video in the GameVibe dataset is $1$ minute long, with every video annotated by the same set of $5$ human annotators. All $30$ games in the dataset have $4$ play sessions. The video clips were selected to contain a maximum of $15$ seconds of non-gameplay content such as pause menus and cut scenes, and were sampled at $30$ hertz with a resolution of $1280\times 720$ for modern titles and $541\times 650$ for older titles. Annotations were collected using the PAGAN annotation platform [44] and the RankTrace annotation tool [37] (see Fig. 3), with the videos presented to participants in random order to minimize habituation and ordering effects. In RankTrace, participants are exposed to stimuli and annotate in real-time by scrolling up or down on a mouse wheel in an unbounded manner to indicate increases and decreases of their labelled state, in this case viewer engagement. Participants of GameVibe were given the following definition of engagement prior to starting their annotation task:
A high level of engagement is associated with a feeling of tension, excitement, and readiness. A low level of engagement is associated with boredom, low interest, and disassociation with the game.
After a qualitative analysis of the dataset, we select $20$ games from the GameVibe corpus to form GameVibe-LLM (see Table I). We discard 10 games that feature third-person segments, mix footage of menus and gameplay, have large mobile UI overlay, or include poor footage. We select one out of four play sessions for each game to generate Few-Shot examples in the final experiments. We test performance on the three remaining play sessions for every game. To be able to fairly compare the performance of different setups, we exclude sessions withheld for the Few-Shot examples from all experiments.
III-B Engagement Data Pre-Processing
Our data preprocessing method closely follows common practices in affective computing and methods introduced in previous studies with GameVibe [56, 55]. Thus, each annotation trace was resampled into $3$-second non-overlapping time windows using simple averaging. The videos were sampled at a similar rate to align the stimuli to the engagement traces provided by the participants. We consider the task a pairwise Preference Learning problem and aim to capture the dynamic change between frames. These traces were then processed into discrete ordinal signals by comparing pairs of consecutive time windows to determine whether engagement increased or decreased between the two time windows. We discard the first comparison in each play session (frame 0 to frame 1) because the very first frame of the videos might lack necessary context. This yields $18$ comparisons per play session—with $4$ play sessions per game, across $20$ games—totalling $1,440$ comparisons.
IV Methodology
In this section we detail our chosen algorithms and the different prompting strategies we employ throughout our experiments. In the presented studies we evaluate the capacity of LLMs to correctly evaluate changes of engagement in gameplay videos. In particular we picked LLaVA and GPT-4o as our base LLMs under investigation (see Section IV-B). In all reported experiments the downstream task of the employed LLM is to label a change in engagement (increase or decrease) given two consecutive frames of a video. We evaluate the algorithm’s performance against the human labelled engagement data of GameVibe that we treat as our ground truth.
To explore how different experimental setups affect LLM engagement predictability, we ran experiments both with Multimodal and Text Descriptions. Figure 2 illustrates the overall strategy and the different experimental settings employed. In the Multimodal Input setting, the input for the algorithm is one or two images accompanied by a text-prompt describing the task. We detail the format of the multimodal input in Section IV-C. In the Text Descriptions setting, instead, we provide text-based descriptions of two video frames as part of the text prompt. We describe the format of the text input in Section IV-D. Finally, we also study few-shot prompting, using multimodal input and we detail this process in Section IV-E along with our general prompting strategy.
IV-A Baseline Algorithms
As a point of comparison to our employed LLMs, we test two baseline algorithms on the GameVibe-LLM corpus. Both baselines were trained using 4-fold cross-validation—i.e. leaving one session out—and using early stopping based on prediction accuracy in the validation set. Both models are trained via cross entropy loss and the Adam optimizer.
For the first baseline, we extracted visual embeddings from frames in the corpus using the pre-trained DinoV2 “gigantic” [53] model, resulting in each image being compressed into a 1536 element vector. Prior to passing images through the model, each channel was normalized using the mean and standard deviations from the ImageNet dataset [21], which was the training data for the backbone models. We then trained a simple MLP classifier to map these embeddings to the engagement labels. The network consists of 2 layers of 1024 and 512 neurons, respectively, and was trained for a maximum of 300 epochs.
Second, we test a convolutional neural network (CNN) classifier using the raw frames as input. The RGB frames were normalized and resized to $(128\times 128)$. The CNN consists of four convolutional blocks, each with a kernel size of 3, followed by two fully connected layers of size $25,088$ and 512, respectively, with a dropout layer (rate of $0.5$) in between.
IV-B Employed LLMs
We employ a selection of different LLMs to provide a comprehensive look on the general capabilites of these models. We sample popular models with multi-modal (multi-image) capabilities from theHuggingface [^5] https://huggingface.co/ open source repository. We focus on models with most downloads and select the following LLMs: Phi 3.5-Vision; Gemma 3; InternVL2.5; LLaVA-Onevision; Qwen2.5-VL; and GPT-4o. The selected open-source LLMs fall between $4$ and $8$ billion parameters. While the models can be considered only ”medium“ in size, their popularity in the community attests their widespread use. We selected GPT-4o as a benchmark for large stat-of-the-art closed-source LLMs. The parameter size of GPT-4o has not been publicly disclosed.
IV-B1 Phi-3.5 Vision
Phi-3.5 Vision (Phi 3.5) is a compact multimodal model from Microsoft, extending the Phi-3 series multimodal capabilities [1]. The architecture consists of a CLIP-based ViT [58] and transformer decoder. To support different resolutions, the system divides images to multiple blocks and concatenates them. The text and image tokens are interleaved in the input space of the LLM backbone. The model is pre-trained to predict text output, then further tuned using both Supervised Fine-Tuning and Direct Preference Optimization [60]. Despite its smaller size ($4.15$ billion parameters ), Phi 3.5 is designed for high efficiency and instruction-following performance, trained on both curated synthetic and human-annotated data during. The model is optimized for reasoning and captioning tasks and provides competitive accuracy with low resource requirements. For our experiments, we use instruction-tuned variant. [^6] https://huggingface.co/microsoft/Phi-3.5-vision-instruct
IV-B2 Gemma 3
Gemma 3 is a member of the third-generation of Gemma models developed by Google DeepMind [68]. Gemma 3 uses a decoder-only transformer architecture and supports both text and vision modalities. The model’s vision encoder is derived from the SigLIP family [86], using $896\times 896$ pixel fixed-resolution tiles. This vision encoder is shared by all Gemma models regardless of the parameter size of the LLM backbone. We chose the $4.3$ billion parameter instruction-tuned variant for our experiments. [^7] https://huggingface.co/google/gemma-3-4b-it Despite the smaller parameter size, Gemma 3 4b was trained on a large budget of just over $4$ trillion tokens to be able to compete with medium sized LLMs. The final model is fine-tuned using a combination of knowledge distillation via an instruction-tuning teacher network and a wide array of Reinforcement-Learning objectives—including weight averaged reward models, human feedback, and ground-truth rewards.
IV-B3 InternVL2.5
InternVL2.5 belongs to a family of robust open-source models. It was developed by OpenGVLab [15] to match the performance of commercial LLMs—such as GPT-4o [51] and Claude-3.5-Sonnet [4]. The model is based on the InternViT Vision Transformer a small 2-layer MLP projector an an LLM backbone. InternVL2.5 is trained iteratively through Native Multimodal Pre-Training, Supervised Fine-Tuning, and Mixed Preference Optimization (MPO) [76]. InternVL2.5 integrates visual and linguistic signals using a fine-grained alignment strategy and is trained on a large multimodal corpus including real-world image-text pairs and synthetic instruction-following data. The 2.5 version introduces improvements in alignment accuracy and spatial reasoning. We chose the $8$ billion parameters (8b) version with the aforementioned MPO fine-tuning, which offers a balance between scalability and performance and has demonstrated strong results on visual question answering. [^8] https://huggingface.co/OpenGVLab/InternVL2_5-8B-MPO
IV-B4 LLaVA-OneVision
LLaVA-OneVision (LLaVA-OV) is a multimodal model based on the architecture of LLaVA [36]. It uses a SigLIP-based vision encoder [86], a 2-layer MLP projector, and Qwen2 as the LLM backbone. The model is trained in a 3-step process, which includes Language-Image Alignment, High-Quality Knowledge Learning, and finally Visual Instruction Tuning [34]. LLaVA-OV was trained on a wide array of single image, multi-image and video tasks and demonstrates strong multimodal understanding. By incorporating the Qwen2 language model, it benefits from a stronger language prior compared to earlier Vicuna-based [16] versions [36]. The model is instruction-tuned to perform general-purpose visual reasoning tasks. We selected the $8$ billion parameter model [^9] https://huggingface.co/llava-hf/llava-onevision-qwen2-7b-ov-hf based on Qwen2 7b for our experiments.
IV-B5 Qwen2.5-VL
Qwen2.5-VL is a multimodal variant of Qwen2.5, developed by Alibaba Group [5]. Qwen2.5-VL uses a custom Vision Transformer (ViT) combined with a 2-layer Multi Layer Perceptron (MLP), which projects the vision features into that aligns with the text embeddings of the LLM backbone. This method allows Qwen2.5-VL to process multiple images of varying resolution (or videos of varying length). The model is trained for structured multimodal reasoning, using spatial and temporal grounding via Multimodal Rotary Position Embedding—introduced in [74]. We use the 7b version of the instruction-tuned variant, which—despite the name—has $8.29$ billion (8.3b) parameters [^10] https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct for our experiments.
IV-B6 GPT-4o
GPT4 is, at the time of writing, the most recent of a series of Generative Pre-trained Transformer (GPT) models developed by OpenAI. GPT4 is a closed source model. While a technical report about GPT4 has been published [51], the exact architecture and training data is unknown. What is known is that GPT4 uses a transformer architecture for both vision and language tasks, relies on reinforcement learning from human feedback and makes use of rule-based reward models based on hidden policy models and human-written rubrics to steer the algorithm in a direction that is considered “safe” by OpenAI. In this paper we use the GPT-4o (Omni) 2024-08-06 model variant. At the time of writing this is considered the flagship model of OpenAI. Unlike previous iterations, GPT-4o is trained end-to-end to incorporate text, audio, image, and video in both its input and output space [52]. We have selected this model because it is one of the most popular [65], state-of-art, closed-source LLMs as an alternative to the open-source LLaVA. We leverage the Open AI API [^11] https://platform.openai.com/ for all reported experiments with GPT-4o.
IV-C Multimodal Input
In our Multimodal experiments we feed the models with both visual input and a text prompt. We extract frames from the GameVibe dataset at a given interval. Despite our selected models capable of handling varying resolutions, for consistency, each frame is cropped to a square and downsized to $512\times 512$. We chose this size because GPT-4o processes images in $512\times 512$ pixel tiles with a maximum image size of $2048\times 768$ [^12] See more information in the OpenAI cookbook:
https://platform.openai.com/docs/guides/vision
. In our Multimodal One-Shot experiments we provide two separate images (consecutive frames) to the model. In out Multimodal Few-Shot experiments, we generate the positive and negative examples in a similar way, using two images and a text prompt as input. For the downstream task of engagement evaluation we provide $6$ images in total for every query (this includes both examples and the final query).
IV-D Text Input
In our experiments using Text Descriptions we feed the models with text descriptions of two video frames as part of the prompt. We obtain these descriptions using the same LLM we use to generate the engagement evaluation. The image input of this task is formatted in the same way as for the Multimodal One-Shot experiments. Contrary to these experiments, however, here we use these images one-by-one and generate descriptions in two different ways. We call these Basic and Advanced Descriptions based on the amount of context given to the model. For the former, we instruct the model to give a brief description, capturing only essential details without subjective commentary based on the setting and layout, enemies, and player action. For obtaining Advanced Descriptions, we instruct the model to also take player engagement into account and generate a description that captures how it might engage the player or viewer. An illustration this process is given in the Appendix; see Figs. LABEL:annex:text-basic and LABEL:annex:text-advanced respectively. For the engagement prediction task, we feed these descriptions to the models in pairs as part of their text prompt. An example of this prompting strategy and the output it produces is given in the Appendix (see Fig. LABEL:annex:text-prediction).
IV-E Prompting Methods
All prompting strategies we use for the engagement evaluation task follow a Chain-of-Thought (CoT) paradigm [78, 90]. We ask the models to provide a comparison between the given input frames, reasoning its analysis of engagement, and finally offering a one-word decision (i.e., engagement increase or decrease). Additionally, for the Multimodal experiments we also generate a description of the visual input before the comparison. In the Multimodal experiments the task of the model is to explicitly output increasing and decreasing labels. We instruct the model to output its answers in a JSON format, which we parse and extract the final decision from; see also Fig. LABEL:annex:oneshot given in the Appendix. In the experiments with Text Descriptions the decision is to pick the most engaging frame (see Fig. LABEL:annex:text-prediction given in the Appendix).
For our Multimodal Few-Shot experiments, we generate artificial reasoning samples for a positive and negative example for each task. We use the same CoT prompt for this process as for the One-Shot experiments. We will call this prompt “CoT prompt” in the remainder of this section. To generate these samples we take the following steps (see also Multimodal Few-Shot in the middle of Fig 2):
  1. 1.
    We take a random example from the same game as presented in the task from an unseen play session.
  2. 2.
    We use the same CoT prompt as for the final engagement evaluation task but modify the prompt leaving only the correct option for the decision.
  3. 3.
    We amend the prompt with the correct evaluation based on the ground truth (see Ground Truth Engagement on Fig. 2).
  4. 4.
    We add a Reasoning Prompt to instruct the model to provide reasoning for the ground truth evaluation.
By removing incorrect options but using the same CoT prompt when generating positive and negative examples, we ensure that the algorithm’s output is formatted the same way as for the downstream task, including the description, comparison, reasoning, and decision. We use these outputs to construct an artificial history of positive and negative examples, which are added to the final prompt for the engagement evaluation task. For this final step we provide the CoT prompt with the example images as a question, and the example output as an answer; then finally we provide a set of unseen images with the CoT prompt and instruct the LLM to evaluate engagement the same way it would for a one-shot experiment. Figures LABEL:annex:fewshot-examples and LABEL:annex:fewshot-prediction given in the Appendix detail the process starting from example generation all the way to engagement prediction.
V Results
Figure 4: Application of the temporal shift ($\Delta t$) hyperparameter to the ground truth. The top red bar (Vision Input) shows an example of individual frames extracted from the gameplay video at a 3-second interval. The bottom green bar (Ground Truth) shows a $\Delta t$ of 2-2 seconds, which means that each window aggregates information 22 seconds before and 11 second after the corresponding video frame.
Figure 5: Sensitivity analysis across hyperparameters $\Delta t$ and $\theta$. The table presents $\Delta A$ values (relative gain in accuracy). $\Delta t$ is the relative shift of the time window to the frame, and $\theta$ is the binary threshold for the split criterion (i.e., increasing or decreasing engagement). The last column shows average $\Delta A$ across all games. The last two rows in each experimental setup ($\Delta t$ and $\theta$ combination) show the MLP and CNN baselines.
This section presents the main results of the experiments performed as follows. In Section V-A we outline the setup of the experiments reported and in Section V-B we discuss our exploratory findings. In Section V-C we examine LLM performance across different input modalities for the engagement evaluation task. Section V-D presents the results of our few-shot prompting experiments, and finally Section V-E takes qualitative lens in our attempt to explain and justify our core findings.
V-A Experimental Setup
We compare the engagement labels generated by LLMs to an engagement ground truth calculated from 3-second time windows of GameVibe annotation traces as outlined in Section III-B. We introduce and vary two hyperparameters in this process:
  1. 1.
    A temporal shift compared to the observed video frame ($\Delta t$). This is similar to what the literature often refers to as input lag [45]. While this correction is generally used to account for reaction time, here we use it to control the temporal difference between the observed frames and the ground truth (see Fig. 4).
  2. 2.
    A preference threshold ($\theta$), taking values between 0 and $1$, that determines whether a difference between the ground truth value of two consecutive time windows is considered a change (increase or decrease) in engagement; e.g. $\theta=0.05$ considers windows which have a difference of more than $5\%$ when evaluating engagement change. We discard labels that vary less than $\theta$, under the assumption that the change—and therefore the decision—is negligible.
While the evaluation follows an ordinal paradigm, we formulate the downstream task of LLMs as binary classification, and ask our models to predict the increase or decrease of perceived engagement between two frames of consecutive time-windows. We discard predictions which could not be interpreted when either for the following occurs: a) the algorithm predicts no change in engagement, b) the LLM generates outputs we could not parse, or c) the model is not able to provide an output. [^13] Good examples of these cases are models not following formatting instructions; providing other answers than “increase” and “decrease” (e.g. “same”); outputting repeating tokens; and—mostly in case of GPT-4o—refusing to provide analysis for unknown reasons, e.g.: “I’m unable to analyse the content of these images. If you can describe the frames, I can help evaluate the change in engagement”.
Based on how the intensity of the action changes in the footage, the class distribution of increasing and decreasing labels can vary between games. At the lower end of the class imbalance a naive Zero Rule Classifier (ZeroR), always predicting the majority class, would have $52\%$ accuracy (e.g. in Far Cry (2004), Blitz Brigade (2013) and Fortnite (2018)) and at the higher end it would have between $70\%$ to $78\%$ (e.g $78\%$ in Counter Strike 1.6 (2003), $73\%$ in Heretic (1994), and $70\%$ in Corridor 7 (1994). To be able to meaningfully compare performance across different games and instructions, we report accuracy gain over a ZeroR baseline [2, 80]—instead of the accuracy values per se—as follows:
$\Delta A=\frac{A_{m}-A_{0}}{A_{0}}$ (1)
where $\Delta A$ is the relative gain in accuracy; $A_{m}$ is the accuracy of the given model; and $A_{0}$ is the accuracy of a ZeroR classifier for a given game. We use the $\Delta A$ measure of performance in all reported experiments in this paper.
V-B Multimodal One-Shot Experiments and Sensitivity Analysis
We experiment with the temporal shift $\Delta t\in\{0,-0.5,-1,-1.5,-2,-2.5,-3\}$ and preference threshold $\theta\in\{0,0.01,0.05,0.1\}$ parameters—introduced in the previous section. Because we experiment on $7$ different algorithms including the MLP and CNN Baseline the combinations of these parameters result in $196$ experimental setups for each game. Due to space considerations we only present the best performing subset of these hyperparameters ($\Delta t\in\{-1,-2\}$ and $\theta\in\{0.01,0.05\}$). We run these experiments with the Multimodal One-Shot strategy as described in Section IV-C. We chose this setup for the initial parameter tuning because this is the most straightforward setup.
Figure 6 presents the $\Delta A$ performance across two $\Delta t$ and $\theta$ values. We can observe that larger $\Delta t$ and $\theta$ values tend to yield higher performance; and it also appears that model size has a high impact on $\Delta A$. Across all experimental setups, we see Phi 3.5 with 4.15b parameters performing the worst of the open source models, while the best models have 8b-8.3b parameters (LLaVA-OV and Qwen2.5 VL, respectively). To measure the significance of the difference between the models and experimental setups we employ repeated paired Student’s t-Tests. We apply the Benjamini-Yekutieli Method [10] to correct our p-values for multiple observations and consider a significance level of $\alpha<0.05$. We make three types of comparisons: 1) we compare models within each experimental configuration ($\Delta t$ and $\theta$ combinations); 2) we compare different experimental configurations within each model; and 3) we compare performance on individual games across all LLMs (but notbaseline models) within each experimental configuration. In the first two cases, our degree of freedom is $19$, while in the last case it is $5$. We make $28$, 66, and $190$ comparisons per test respectively.
Despite GPT-4o outperforming all other algorithms at every turn, after the correction we find no significant difference between models or different experimental setups. Nevertheless the best performing hyperparameter set is $\Delta t=-2$ and $\theta=0.05$ in terms of average performance. Focusing on this setup, we observe that the best overall performance can be attributed to GPT-4o with $1\%$ $\Delta A$; while the worst LLM is Phi 3.5 with an overall $13\%$ $\Delta A$. When it comes to the comparison against the baselines, all LLMs outperform the MLP Baseline across all experimental setups; and generally provide comparable performance to the CNN Baseline.
Interestingly, while GPT-4o provides the best performance, we can see comparable results with other models and configurations on single games. Most notably Doom (1993) is remarkably easy to predict for all models, including the MLP Baseline, while Counter Strike 1.6 (2003) is the hardest. The difference in performance is significant, $t(5)=27.2$, $p=0.001$ (corrected). In fact, Doom is significantly easier to predict than Corridor 7 (1994), $t(5)=9.6$, $p=0.023$; HROT (2023), $t(5)=8.7$, $p=0.026$; and Team Fortress 2 (2007), $t(5)=6.9$, $p=0.048$ —and it is not the only one. Wolfram (2012) yields significantly better results than Corridor 7, $t(5)=13.2$, $p=0.012$; Medal of Honor (2010), $t(5)=13.8$, $p=0.012$; Counter Strike 1.6, $t(5)=10.1$, $p=0.023$; PUBG (2018), $t(5)=9.8$, $p=0.023$; and HROT, $t(5)=9.3$, $p=0.023$; and Blitz Brigade (2013) is significantly better than Corridor 7, $t(5)=11.6$, $p=0.019$; Counter Strike 1.6, $t(5)=22.5$, $p=0.002$; and HROT, $t(5)=8.1$, $p=0.034$. In contrast to Wolfram, Doom, and Blitz Brigade—with an average $\Delta A$ of $19\%$, $18\%$, and $7\%$ respectively—the worst performing games are HROT, Corridor 7, and Counter Strike 1.6—with an average $\Delta A$ of $28\%$, $41\%$, and $50\%$, respectively across all LLMs in the $\Delta t=-2$ and $\theta=0.05$ setting.
Notably, in games which are easier to predict the proportion of the ground truth majority class is lower—Wolfram ($57\%$), Doom ($53\%$), and Blitz Brigade ($51\%$)—while games where the change in engagement is harder to predict have a higher naive baselineHROT ($61\%$), Corridor 7 ($71\%$), and Counter Strike 1.6 ($78\%$). This indicates that engagement prediction is easier in where the provided frames capture a more dynamic gameplay footage and have a more uniform distribution of increasing vs. decreasing engagement labels.
Considering the overall performance of LLM engagement prediction across games, we fix our parameters for processing the ground truth at $\Delta t=-2s$ and $\theta=0.05$ for the remaining experiments presented in this paper.
V-C Engagement Prediction Based on Text Descriptions
Figure 6: Model performance on Text Descriptions across Basic and Advanced descriptions, and the best Multimodal One-Shot results identified in Section V-B. The table presents $\Delta A$ values (relative gain in accuracy). The last column show the average $\Delta A$ across all games. The last two rows show the MLP and CNN Baselines.
In this section we examine the impact of text-based vs. multimodal prompting strategies on LLM performance. While in the former case we provide solely a text prompt to the model, in the latter case we feed both a text prompt and a corresponding image. Because the performance of LLMs can be affected even by small prompt variations [41], we experiment with both Basic and Advanced prompts. The prompting procedure for the text-based experiments are detailed in Section IV-D. Figure 5 shows the $\Delta A$ performance of Text Descriptions experiments compared to the best Multimodal One-Shot models discussed in the previous section.
In this section our analysis focuses on the Text Descriptions compared to the Multimodal One-Shot results presented in the previous section. This focus on text allows us to compare the Text Descriptions method to a simple multimodal approach across different models. Our hypothesis is that the strategy of generating text-descriptions of frames first and then using these descriptions as part of the Text Descriptions will improve model performance, because it essentially encodes the images in terms of action and player involvement. We assume that using this type of Text Descriptions will present a better representation by discarding surface-level differences between frames and emphasising the structural differences.
We make four types of comparison: 1) we compare models within the Basic and Advanced Text Description tasks; 2) we compare between the downstream tasks of LLMs (Basic and Advanced Text Description, and Multimodal One-Shot) and the Baseline performance within each model; 3) we compare performance on individual games across all LLMs within the Basic and Advanced Text Description tasks; and finally 4) we compare performance on individual games between the the downstream tasks of LLMs (Basic and Advanced Text Description, and Multimodal One-Shot). In the first two cases, our degree of freedom is 19, while in the last two cases it is 5. We make 15, 17, 190, and 60 comparisons per test, respectively.
In general, the overall results of the experiments on Text Descriptions produce similar results to the Multimodal One-Shot setup. While we can observe that the overall performance is better on Advanced Text Descriptions than on Basic Text Descriptions; and that Multimodal One-Shot outperforms both—when we compare models both within tasks and between tasks, no significant results are observed. The same is true when comparing them to the MPL and CNN Baselines. However, several significant differences emerge when comparing individual game performance across all LLMs within the Basic and Advanced Text Description tasks. Most notably, Wolfram significantly outperforms PUBG, $t(5)=17.8$, $p=0.011$ and Counter Strike 1.6, $t(5)=14.6$, $p=0.015$; and Fortnite outperforms Counter Strike 1.6, $t(5)=11.5$, $p=0.032$ in the Advanced Text task. In the Basic Text task, Wolfram outperforms Medal of Honor, $t(5)=8.8$, $p=0.031$; Corridor 7, $t(5)=9.6$, $p=0.029$; Counter Strike 1.6, $t(5)=9.7$, $p=0.029$; and Borderlands 3, $t(5)=10.2$, $p=0.029$. Doom outperforms Heretic, $t(5)=11.1$, $p=0.028$; Counter Strike 1.6, $t(5)=11.2$, $p=0.028$; and Void Bastards, $t(5)=9.2$, $p=0.029$. Finally, both FarCry and Blitz Brigade outperforms Counter Strike 1.6, $t(5)=14.0$, $p=0.023$ and $t(5)=13.3$, $p=0.023$; and Corridor 7, $t(5)=9.1$, $p=0.029$ and $t(5)=9.5$, $p=0.029$ respectively. Finally, comparing task performance within individual games, we find that the Multimodal One-Shot task significantly outperforms Advanced Text Description on Counter Strike 1.6, $t(5)=5.0$, $p=0.023$; while both Basic and Advanced text tasks outperform Multimodal on Borderlands 3 (2019), $t(5)=3.5$, $p=0.047$ and $t(5)=4.0$, $p=0.047$ respectively.
Similarly to the Multimodal One-Shot experiments, Wolfram, Doom, and Blitz Brigade are the easiest games to predict; and Counter Strike 1.6, Heretic, and Corridor 7 are the hardest to predict. While there are some standout performances that we can observe on 5—for example Phi 3.5 is comparable to GPT-4o on Wolfram but performs poorly on Doom in the Advanced Text Description task compared to every other model—there is no apparent overarching pattern we can analyse. It also seems that any performance outliers can mostly be explained through the particularities of the data and the chosen algorithm.
We started this section with a hypothesis that using text-descriptions of frames would improve predictive capacity of LLMs compared to multimodal inputs. We believed this would be the case because the generation of text descriptions would act as a type of game-agnostic encoding, putting more emphasis on the layout and action of frames. The results presented here indicate that this is not the case. In general, obtained results show no significant differences between the Text Descriptions and Multimodal One-Shot setups. Generating text-descriptions first and using text-only input cannot provide a better encoding than simpler multimodal approaches for this task. Because Multimodal LLMs were trained to encode images and text into a shared embedding space [58] the extra “image to text” step is unnecessary.
V-D Multimodal Few-Shot Prompting
Figure 7: Results of the Multimodal Few-Shot experiments across all models compared to the best Multimodal One-Shot results from Section V-B. The last column show the average $\Delta A$ across all games. The last two rows show the MLP and CNN Baselines.
In this section we present experiments using Multimodal Few-Shot strategy (see Fig. 2 and Section IV for more details on this approach). As there is no significant difference between the Multimodal One-Shot experiments and those that are using Text Descriptions, here we only focus on comparisons between the Multimodal setups. Figure 7 presents the results of our Multimodal Few-Shot experiments compared to the best Multimodal One-Shot results presented in Section V-B. We can observe an overall improvement of performance in the Few-Shot setup.
Similarly to the experiments based on Text Descriptions, we make the following comparisons: 1) Few-Shot experiments against each other; 2) Few-Shot experiments against One-Shot and Baseline experiments within the same model; 3) performance on individual games across all Few-Shot experiments; and finally 4) performance on individual games between Few-Shot and One-Shot setups. In the first two cases, our degree of freedom is 19, while in the last two cases it is 5. We make 15, 8, 190, and 20 comparisons per test, respectively.
When we compare the average performance of models on the Few-Shot setup, we observe that GPT-4o outperforms all other models. However, when we compare the results across games, this outstanding performance is only significant when comparing it to InternVL2.5, $t(19)=4.4$, $p=0.015$. Comparing models between tasks (Few-Shot, One-Shot, and Baselines), the performance differences between Few-Shot and One-Shot tasks are not significant. Nevertheless, GPT-4o significantly outperforms both the MLP and CNN Baselines, $t(19)=4.3$, $p=0.005$ and $t(5)=4.2$, $p=0.005$ respectively. Interestingly, while we can see that games which have been easy or hard to predict in the One-Shot setting are consistent in the Few-Shot setting, there is no significant difference between games. Nevertheless, we can observe that once again, Wolfram, Blitz Brigade, and Doom define the easiest downstream tasks; and Counter Strike 1.6, Medal of Honor, and Corridor 7 are among the hardest to predict. Comparing task performance within single games, we can observe improvement when going from One-Shot to Few-Shot prompting. This is especially true in the otherwise difficult-to-predict games: Corridor 7 ($17\%$); HROT ($13\%$); and Counter Strike 1.6 ($12\%$). Nevertheless, the only game where the performance difference across all models is significant is CS:GO - Dust2 (2012), $t(5)=2.9$, p=0.034p=0.034—where we see a $7\%$ drop.
While there are few significant differences between the One-Shot and Few-Shot experiments, there is a clear pattern of overall improvement when it comes to GPT-4o, Gemma 3, and Qwen2.5-VL. Most notably GPT-4o achieves an average $6\%$ $\Delta A$ and Qwen2.5-VL shows $59\%$ $\Delta A$ on Wolfram—the absolute best result across all experiments. However, similarly to previous experiments, there are also negative outliers. Performance in Doom is dropped by $10\%$; and in case of LLaVA-OV we can see an overall $5\%$ drop between One-Shot and Few-Shot setups. Because there is no clear pattern between models, differences must come down to the specifics of the data and the architectures of individual models.
V-E Qualitative Analysis
In this section we outline the reasons for the observed poor performance of the tested LLMs and analyse why certain games are easier to predict. For our analysis we are looking at the highest performing model, the GPT-4o with Multimodal Few-Shot prompting. Employing this model we list $5$ games where the $\Delta A$ exceeds $25\%$: Doom, Wolfram, PlayerUnknown’s Battlegrounds (PUBG) (2018), Apex Legends, and Borderlands 3 (2019). Conversely, the five games, where the performance was well-below the baseline are as follows: Corridor 7, Heretic, Counter Strike 1.6, Medal of Honor (2010), and HROT; see Fig. 8.
Figure 8: The 5 best and worst performing games in terms of $\Delta A$ (relative gain in accuracy) using Multimodal Input - 2 Images with GPT-4o Few-Shot prompting. Best games from left to right: a) Doom, b) Wolfram, c) PlayerUnknown’s Battlegrounds (PUBG), d) Apex Legends, and e) Borderlands 3. Worst games from left to right: f) Corridor 7, g) Heretic, h) Counter Strike 1.6, i) Medal of Honor 2010, j) HROT
A qualitative analysis of the games where LLMs perform best (vs. those where they perform worst) reveals some possible underlying reasons that could influence these models. The five games where LLMs perform best are fast paced, with short bursts of action separated by similarly short navigation sequences. The game scenes are well-lit or stylized in a way that is easy to read. In contrast, the five games where LLMs fail to assign engagement labels feature repetitive sections of navigation with limited gameplaying action such as shooting, reloading, collecting items, or dodging fire. These games also tend to feature dark backgrounds and enemies with silhouettes that are difficult to distinguish, or they take place in drab environments where the ground, background, and often non-player characters blend together.
Figure 9: Similar frames from Counter Strike variants: Counter Strike 1.6, CS:GO - Dust2, and CS:GO - Office (left to right)
A representative example that highlights these performance differences are the Counter Strike game variants existent in the dataset; see Fig. 9. Compared to the best performance of the Multimodal Few-Shot GPT-4o on Counter Strike 1.6 ($33\%$ worse than baseline), the same model on CS:GO - Dust2 has a performance comparable to baseline levels. Even though these two games use essentially the same level, the visuals of CS:GO - Dust2 are much clearer; in Counter Strike 1.6 the background and foreground are harder to separate visually. In CS:GO - Office—where the visuals are arguably even more readable—the model showcases much higher predictive capacity (i.e. $19\%$ higher than the baseline).
Another way to explain the fluctuation in LLM performance is the familiarity of the model with the games per se. We observe that more popular games (such as Counter Strike, Apex Legends, and PUBG, with a peak viewership [^14] Peak viewership refers to the historically highest number of concurrent viewers watching a stream. It is indicative of the maximum audience size. of $1,914,861$, $674,070$, and $597,663$, respectively on Twitch [^15] Numbers retrieved from https://twitchtracker.com/, January 2025.) yield generally better engagement predictions compared to less popular games (such as HROT, Heretic, and Corridor 7, with a peak viewership of $24,721$, $2,280$, and $195$ on Twitch15), although we should be careful with naive over-generalizations from these findings. For one, Counter Strike 1.6 is a variant of a very popular game with a peak viewership of $125,378$ in itself, but the models struggle with correctly evaluating the change in engagement—at least in the GameVibe dataset. While it is possible that the training data of GPT-4o contains images from more popular games, attempting to verify this by reconstructing parts of the GPT-4o training data is out of the scope of this paper.
VI Discussion
The evaluation experiments presented in this paper are the first of its kind for LLM-based engagement prediction in games. While collectively we tried $4,840$ combinations of experimental settings—varying the model type, prompting strategy, input type, game, and ground truth—there are still many aspects that we did not explore in this initial study. We argue, however, that we set out to lay ground work for future research by approaching the problem of automating gameplay annotation in a relatively straightforward way.
VI-A Limitations
The study presented here constitutes a preliminary look into how out-of-box LLMs can solve affective computing tasks relying on their knowledge priors. Naturally, the results of such a study face multiple limitations.
VI-A1 Static Images
We experimented with several models and prompting strategies but kept the granularity of the vision input constant. We focused on single frames sampled at a 3-second interval because the low resolution input is less costly and more readily available than high frames-per-second videos. Nevertheless, there is a high likelihood that some of the context of the action is lost during this processing. This is a core limitation of this initial study. Although we briefly experimented with different time intervals (i.e. between 1 and 5 seconds), simply increasing the sampling rate did not yield a performance increase. It is likely, however, that by either providing more frames per query or using video input directly would lead to a significant performance improvement that remains to be tested in future studies.
VI-A2 No Fine-Tuning
As our focus was to evaluate the state of LLMs relying solely on their own knowledge priors, we did not employ any form of fine-tuning. The main motivation for this decision is that we are interested in the zero-shot capabilities of available LLMs and their limits using prompt-engineering. While we can expect models with specialised fine-tuning to outperform general models, this is not always the case. Large commercial LLMs using few-shot prompting have shown greater flexibility and superior performance compared to pre-trained models that have been fine-tuned on specialised datasets in certain domains [77]. As full-parameter fine-tuning is prohibitively costly, it is worthwhile investigating other avenues of making use of these models. Here we demonstrated prompt-engineering including a few-shot prompting strategy, but future studies should investigate efficient parameter tuning, such as low-rank adaptation [49, 54].
VI-A3 Limited Domain
We showcase our results on $20$ different games with varying aesthetics and play-styles, however, the experiments focus only on the first-person shooter genre. Even though this is a popular videogame genre—with countless examples out there, we show that positive results generalise poorly even within this limited domain. Our results show that clear visuals are key to the models’ success. This holds true across different experiments and architectures. These results show the need for generalised game representation beyond simple pixels on the screen [71].
VI-B Future Work
While video input could feed more information to the LLM, the context of the query could also be augmented and then provided to the LLM, thereby improving its predictive capacity. By implementing a memory mechanism [89], we could potentially store and recall the temporal context of the play session. Similarly, we could provide more context on the necessary domain knowledge for the task by implementing retrieval-augmented generation [33], where we could feed more information on the game or downstream task. We plan to pursue these avenues in our future studies in our effort to further investigate how more contextual information impacts the performance of LLMs towards fully autonomous engagement annotation.
Generating subjective labels is a relatively open field with a lot of unanswered questions. Naturally, the exploration should be extended to other datasets. While this study focusses on engagement, there are other subjective aspects of both player and viewer experience that could be evaluated further. A natural step forward would be to make use of a diverse set of affective corpora, focusing, for instance, on affect prediction across videogame datasets [46], but also architectural spaces [79] and movie corpora [27]. As discussed above, the current evaluation of LLMs—even though multimodal—considers a predetermined number of modalities: text and images. As we move forward and more multimodal architectures become widely adopted, the research into utilizing LLMs for autonomous affect annotation could encompass different modalities from images, through video, to audio. When it comes to interactive media such as games, user behavioural data could also be included [62] providing a richer context to the models.
VII Conclusion
This paper explored a novel application of LLMs for autonomously annotating the continuous experience of viewers when consuming videos of first-person shooter video games from the GameVibe corpus. We conducted an in-depth analysis comparing multiple open-source foundation models—comparing them to Open-AI’s GPT—and evaluated their performance across different input modalities (i.e. multimodal, text-based) and prompting strategies (i.e. one-shot, few-shot). While results obtained are far from what one could consider as good performance, all LLMs examined surpass the performance of our traditional ML baseline methods across all experiments. This shows the potential of LLMs in otherwise complex affect modelling tasks. Our findings confirm that model size and prompting strategy have a critical impact on model performance. The LLMs presented here demonstrate promising capabilities on certain game elicitors—although their overall performance only marginally surpasses the baseline. Perhaps unsurprisingly, the games where LLMs are successful at predicting the continuos change in engagement are popular games with easy-to-read graphical styles and concise gameplay. The gap in performance on these games compared to more challenging elicitors shows that while LLMs have potential, there is still a long road ahead towards automated continuous affect labelling using these type of foundation models.
As LLMs continue to scale and evolve, we believe their ability to capture subjective experiences will drastically improve—especially when incorporating richer multimodal inputs such as video, audio, and physiological signals. The annotation capacities of such LLM-based foundation affect models extends well beyond the domain of games to video-based general affect modelling, and human-computer interaction at large. Based on the results presented here, we believe future work could leverage LLMs as flexible and scalable annotators in a wide variety of dynamic and real-world settings.
Acknowledgment
This research has been supported by the Foundation.AI project funded by the Maltese Ministry for Education, Sport, Youth, Research and Innovation (MEYR), in consultation with the University of Malta (UM) and the Malta Chamber of Commerce, Enterprise and Industry (MCCEI).